OpenStudy (anonymous):

verify sin^4x-cos^4x=1-2sin^2x

OpenStudy (ipwnbunnies):

Start by factoring the left side. Use the difference of squares idea, like in quadratic equations.

OpenStudy (anonymous):

like this (sin^2x+cos^2x)(sin^2+cos^2x)?

OpenStudy (solomonzelman):

Close, it's + and -.

OpenStudy (anonymous):

and what do I do afterwards?

OpenStudy (anonymous):

do i turn sin^2x+cos^2x into a 1?

OpenStudy (solomonzelman):

yes. good thinking !

OpenStudy (anonymous):

ok so i turned cos^2x into 1-2sin^2x but i'm left with -1-2sin^2x

OpenStudy (solomonzelman):

sin^4x-cos^4x=1-2sin^2x (sin^2x+cos^2x)(sin^2x-cos^2x)=1-2sin^2x (1)(sin^2x-cos^2x)=1-2sin^2x sin^2x-cos^2x=1-2sin^2x -cos^2x=1-3sin^2x This is what you are left with so far...

OpenStudy (anonymous):

yeah so -1-2sin^2x =/= 1-2sin^2x , what do I do now?

OpenStudy (solomonzelman):

where do you get that from ?

OpenStudy (solomonzelman):

@Enlightened , can you work both sides at the same time, or you only can work with 1 side ?

OpenStudy (anonymous):

1 side? cos^2x = 1-2sin^2x but how do I get rid of the negative?

OpenStudy (solomonzelman):

sin^2x-cos^2x=1-2sin^2x +sin^2x +sin^2x -cos^2x=1-sin^2x THE COSINE IS NEGATIVE.

OpenStudy (solomonzelman):

-1+sin^2x=1-sin^2x is certainly NOT always the case. I hate verifying, without being able to use both sides....

OpenStudy (anonymous):

well it's the side I have to prove, i don't think you can use both sides.